# L1: Controlling How Long A Reasoning Model Thinks With Reinforcement Learning

### v1

@@ Abstract
Reasoning language models have shown an uncanny ability to improve performance at test-time by “thinking longer”—that is, by generating longer chain-of-thought sequences and hence using more compute. However, the length of their chain-of-thought reasoning is not controllable, making it impossible to allocate test-time compute to achieve a desired level of performance. We introduce Length Controlled Policy Optimization (LCPO), a simple reinforcement learning method that optimizes for accuracy and adherence to user-specified length constraints. We use LCPO to train L1, a reasoning language model that produces outputs satisfying a length constraint given in its prompt. L1’s length control allows for smoothly trading off computational cost and accuracy on a wide range of tasks, and outperforms the state-of-the-art S1 method for length control. Furthermore, we uncover an unexpected short chain-of-thought capability in models trained with LCPO. For instance, our 1.5B L1 model surpasses GPT-4o at equal reasoning lengths. Overall, LCPO enables precise control over reasoning length, allowing for fine-grained allocation of test-time compute and accuracy. [^1] Code and models released at https://cmu-l3.github.io/l1
@@ 1 Introduction
Beyond improved length control in the standard math reasoning setting, we find that LCPO-trained models generalize surprisingly well to out-of-distribution tasks, including logical reasoning, and general-knowledge benchmarks like MMLU (Hendrycks et al., 2021a). Furthermore, we show that “long-CoT” models trained with LCPO become unexpectedly strong “short-CoT” models: when prompted to generate short reasoning traces, LCPO-trained models outperform their original counterparts by significant margins (up to 10% improvement), even at the same generation length. To the best of our knowledge, for the first time we show that a 1.5B model can match the performance of GPT-4o (OpenAI et al., 2024b) despite using the same token budget. In summary, our contributions are:
@@ 2 Related Work
Length control specifically tailored to reasoning tasks remains relatively unexplored. Recent works, such as those by Arora & Zanette (2025) and Kang et al. (2024), emphasize generating shorter reasoning chains for efficiency, but they do not enable explicit length control or precise alignment with user-specified inference budgets. Another work S1 (Muennighoff et al., 2025) introduces “budget-forcing” by imposing a strict token limit: either truncating output at budget exhaustion or inserting a special token (“Wait”) to request continued generation until reaching the full length budget. Unfortunately, this strategy presents significant practical drawbacks. Abrupt truncation often interrupts reasoning mid-step, negatively impacting model accuracy and user interpretability. Meanwhile, the repetitive usage of special continuation tokens risks rigid and suboptimal reasoning patterns.
In contrast to these prior works, our LCPO is uniquely designed to train reasoning-specialized models for precise and adaptive length control. LCPO uses reinforcement learning so that models learn to dynamically allocate inference compute based on constraints provided in a prompt. As our experiments will demonstrate, our method substantially surpasses previous approaches in precision over length control, and performance at varying length budgets.
@@ 3 Method
$x_{i}^{new}=\text{Concat}\Bigl{(}x_{i},\,\text{``Think for }n_{gold,i}\text{ tokens.''}\Bigr{)},$.” ) ,
@@ 4 Experimental Setup
  • DeepSeek-R1-Distill-Qwen-1.5B: is the SFT version of Qwen-2.5-1.5B-Instruct finetuned on reasoning traces of DeepSeek’s R1 model. For brevity, we refer to this model as DeepSeek-R1-1.5B.
  • DeepScaleR-1.5B-Preview: the original model, evaluated without any length control modifications. For brevity, we refer this model as Agentica-24K.
  • DeepScaleR-1.5B-Preview-4K: a version of Agentic-24K fine-tuned with 4K context length. This is done due to computational constraints of training LCPO with long sequence length (such as 24K used in Agentica-24K). The model therefore serves as a fair comparison to L1. For brevity, we refer to this model as Agentica-4K.
  • S1: (Muennighoff et al., 2025) is a budget-forcing method, which controls reasoning length using simple test-time interventions. We implement this method on top of the Agentica-24K model.
@@
We evaluate our approaches along two dimensions. First, we assess the model’s ability to adhere to the targeted length by reporting the mean deviation between the generated token length $n_{y}$ and the target $n_{gold}$. Second, we evaluate the overall performance (i.e., problem-solving accuracy) when generating responses at different target lengths. In our experiments, target lengths are selected from $\{512,1024,2048,3600\}$ tokens.
@@ 5 Results and Analysis
Figure 2: Performance of L1 under varying token budgets. Pass rate (y-axis) across multiple benchmarks as a function of tokens used (x-axis). L1-Exact and L1-Exact significantly outperform length-controlled baselines (S1) while approaching the performance of unconstrained models.
In this section, we report and analyze the effectiveness of the proposed method (LCPO) across various settings and benchmarks. We evaluate our method’s relative performance, generalization capability on out-of-domain tasks, controllability of length constraints and competitive performance in short Chain-of-Thought (CoT) setups, and examine learned reasoning behaviors.
@@
Figure 2 compares performance of L1-Exact and L1-Max with other baselines across varying generation lengths. Both variants of L1 achieve superior performance across all token budgets while maintaining precise length control. Compared to S1, the only other method specifically designed for length control, L1 shows remarkable improvements, over 100-150% relative and 20-25% absolute performance gains at both 512 and 1024 token budgets. This substantial difference can be attributed to two key factors: (1) L1 intelligently adapts its chain-of-thought to fit within specified length constraints without disrupting the reasoning process, while S1 often truncates mid-reasoning; and (2) L1 is explicitly trained to generate high-quality reasoning chains of varying lengths, effectively distilling reasoning patterns from longer chains to shorter ones.
@@
L1-Exact performs approximately 1% below Agentica-4K, which is the same underlying model as L1, but trained without length constraints. However, this difference is primarily observed in the AIME dataset, where unconstrained models can generate very long chains for complex problems. Additionally, L1-Exact allocates the same token budget to all problems regardless of difficulty, potentially using extra tokens on simpler problems. L1-Max effectively alleviates this challenge, matching the performance of Agentica-4K by optimizing token usage based on problem difficulty while respecting the upper ceiling. In doing so, it outperforms even L1-Exact often by up to 2x fewer tokens. L1-Max is particularly valuable when exact token counts are less desirable than a worst-case compute budget. Finally, the scaling trends suggest that with longer context training, L1 would match or even surpass Agentica-24K’s performance while maintaining a high degree of length control.
@@
Figure 3 confirms that L1 generalizes robustly to new domains: performance consistently scales positively with token budget for OOD general reasoning datasets, approaching or matching Agentica-4K benchmarks despite explicit length control constraints. For GPQA and LSAT, we observe the same linear performance scaling trend as in our primary datasets, with L1 matching Agentica-4K’s performance at comparable token budgets. This generalization is particularly impressive given that L1 was not explicitly trained on these tasks. For MMLU, we see a less pronounced linear scaling relationship ($R^{2}=0.66$), likely because these knowledge-focused questions benefit less from extended reasoning.
@@
We quantitatively assess L1’s ability to follow length constraints across various mathematical reasoning datasets. As shown in Figure 5, our model maintains consistent control across all token budgets (512, 1024, 2048, and 3600 tokens), with observed output lengths usually closely matching the requested lengths. Further, in Figure 5, we show the mean error: ($\frac{E_{x\sim D}[n_{generated}]-n_{gold}}{n_{gold}}$), which captures the average deviation from target lengths across the dataset. The figure demonstrates that mean error is low: close to 3% for all math reasoning datasets. Although OOD datasets exhibit predictably higher errors (20-40%), these remain preferable over uncontrolled prompting. Further Analysis in Appendix A.2 demonstrates that larger errors primarily appear at higher token budgets on tasks like MMLU, where the longer chain of thoughts is mostly unnecessary. Additionally, in Appendix A.1, we show ethat rror can be further reduced significantly with extended RL training.
@@
Table 1: Short CoT performance across math benchmarks. L1-Max surpasses its non-reasoning base model and competes with frontier models under identical short CoT lengths. * Our analysis reveals model is significantly contaminated, See results on AIME2025 for details.
Given L1’s strong performance at lower token budgets, we conducted a focused evaluation comparing it to both its base non-reasoning model (Qwen-2.5-1.5B-Instruct) and significantly larger non-reasoning models (GPT-4o and Llama-3.3-70B) at comparable generation lengths. Table 1 presents these results, showing that L1 consistently outperforms or matches all models across all datasets despite using equivalent token budgets. Further, on average, L1 is 5% better than its non-reasoning counterpart, and even outperforms GPT-4o by 2% on average.
@@
Figure 6: Keyword usage by reasoning pattern at 512 vs. 4096 tokens. L1 allocates more tokens to self-correction and conclusion drawing at higher budgets, reflecting adaptive reasoning strategies.
To understand how L1 changes its reasoning approach across different length constraints, we analyzed how frequently certain reasoning-related terms appear in outputs of different lengths. Specifically, we calculated the normalized occurrence rate of most common reasoning terms in 512-token outputs compared to 4096-token outputs, showing how the model’s reasoning strategies shift when given different length constraints. Figure 6 organizes these keywords into four distinct reasoning patterns: “Self-Correction and Verification,” “Exploration and Alternatives,” “Context Setting,” and “Conclusion Drawing.”
Figure 7: Distribution of tokens for thinking vs. solution across chain-of-thought lengths. L1 maintains a mostly consistent ratio of thinking tokens to solution tokens, indicating efficient scaling of intermediate steps without making final answers overly verbose.
Figure 6 shows that self-correction and verification keywords appear approximately twice as frequently in 4096-token outputs compared to 512-token outputs. Similarly, conclusion-drawing terms increase 2-10x with increased token budget, indicating more thorough solution validation and completion. Interestingly, most exploration-related keywords decrease in relative frequency at higher token counts, with “Alternatively” being a notable exception. Overall, we observe that smaller CoTs have reasoning patterns similar to their longer counterparts, but with changed relative frequencies that favor more self-verification and conclusion drawing in longer chains-of-thought.
@@ 6 Conclusion
In this work, we introduced Length Controlled Policy Optimization (LCPO), a simple yet powerful reinforcement learning-based method enabling adaptive control over the length of reasoning chains in language models. We use LCPO to train L1, a reasoning language model, optimized to produce outputs that adhere to length constraints given in its prompt. LCPO significantly surpasses previous test-time scaling methods, achieving over 100% relative and 20% absolute improvements in mathematical reasoning tasks compared to prior length-control approaches.
Further, we demonstrated that L1 generalizes robustly beyond its training distribution, extending its length-controlled capabilities to out-of-domain tasks. Furthermore, our analysis revealed an intriguing phenomenon: models trained to generate longer reasoning chains become unexpectedly strong short-CoT reasoners, outperforming significantly larger frontier models such as GPT-4o at identical generation lengths. By providing length control using simple prompt, LCPO opens promising avenues toward more efficient, flexible, and scalable reasoning models.

### v2

@@ Abstract
Reasoning language models have shown an uncanny ability to improve performance at test-time by “thinking longer”—that is, by generating longer chain-of-thought sequences and hence using more compute. However, the length of their chain-of-thought reasoning is not controllable, making it impossible to allocate test-time compute to achieve a desired level of performance. We introduce Length Controlled Policy Optimization (LCPO), a simple reinforcement learning method that optimizes for accuracy and adherence to user-specified length constraints. We use LCPO to train L1, a reasoning language model that produces outputs satisfying a length constraint given in its prompt. L1’s length control allows for smoothly trading off computational cost and accuracy on a wide range of tasks, and outperforms the state-of-the-art S1 method for length control. Furthermore, we uncover an unexpected short chain-of-thought capability in models trained with LCPO. Specifically, using LCPO we derive Short Reasoning Models (SRMs), that exhibit similar reasoning patterns as full-length reasoning models, but can generate CoT lengths comparable to non-reasoning models. They demonstrate significant performance gains, for instance, our 1.5B L1 model surpasses GPT-4o at equal reasoning lengths. Overall, LCPO enables precise control over reasoning length, allowing for fine-grained allocation of test-time compute and accuracy.  [^1] Code and models released at https://cmu-l3.github.io/l1
@@ 1 Introduction
Beyond improved length control in the standard math reasoning setting, we find that LCPO-trained models generalize surprisingly well to out-of-distribution tasks, including logical reasoning, and general-knowledge benchmarks like MMLU (Hendrycks et al., 2021a). Furthermore, we show that “long-CoT” models trained with LCPO become unexpectedly strong “short-CoT” models: when prompted to generate short reasoning traces, LCPO-trained models outperform their original counterparts by significant margins (up to 10% improvement), even at the same generation length. We refer these models as Short Reasoning Models (SRMs). To the best of our knowledge, for the first time we show that a 1.5B model can match the performance of GPT-4o (OpenAI et al., 2024b) despite using the same token budget. In summary, our contributions are:
@@ 2 Related Work
Length control specifically tailored to reasoning tasks remains relatively unexplored. Prior works have shown that reasoning models often underthink or overthink, both hurting performance and wasting compute (Wang et al., 2025; Chen et al., 2025). Recent works, such as those by Arora & Zanette (2025) and Kang et al. (2024), emphasize generating shorter reasoning chains for efficiency, but they do not enable explicit length control or precise alignment with user-specified inference budgets. Another work S1 (Muennighoff et al., 2025) introduces “budget-forcing” by imposing a strict token limit: either truncating output at budget exhaustion or inserting a special token (“Wait”) to request continued generation until reaching the full length budget. Unfortunately, this strategy presents significant practical drawbacks. Abrupt truncation often interrupts reasoning mid-step, negatively impacting model accuracy and user interpretability. Meanwhile, the repetitive usage of special continuation tokens risks rigid and suboptimal reasoning patterns.
In contrast to these prior works, LCPO is uniquely designed to train reasoning-specialized models for precise and adaptive length control. LCPO uses reinforcement learning so that models learn to dynamically allocate inference compute based on constraints provided in a prompt. As our experiments will demonstrate, our method substantially surpasses previous approaches in precision over length control, and performance at varying length budgets.
@@ 3 Method
$x_{i}^{new}=\text{Concat}\Bigl(x_{i},\,\text{``Think for }n_{gold,i}\text{ tokens.''}\Bigr),$
@@ 4 Experimental Setup
1.) DeepSeek-R1-Distill-Qwen-1.5B: is the version of Qwen-2.5-1.5B-Instruct finetuned on reasoning traces of DeepSeek’s R1 model. For brevity, we refer it as DeepSeek-R1-1.5B.
2.) DeepScaleR-1.5B-Preview: the original model, evaluated without any length control modifications. For brevity, we refer this model as DeepScaleR-24K.
3.) DeepScaleR-1.5B-Preview-4K: a version of DeepScaleR-24K fine-tuned with 4K context length. This is done due to computational constraints of training LCPO with long sequence length (such as 24K used in DeepScaleR-24K). The model therefore serves as a fair comparison to L1. For brevity, we refer to this model as DeepScaleR-4K.
4.) S1: (Muennighoff et al., 2025) is a budget-forcing method, which controls reasoning length using simple test-time interventions. Concretely, once the maximum token budget is reached, S1 stops further generation and instead inserts “Final Answer” in the prompt to force model to generate the final answer.
@@
We evaluate our approaches along two dimensions. First, we assess the model’s ability to adhere to the targeted length by reporting the mean deviation between the generated token length $n_{y}$ and the target $n_{gold}$. Second, we evaluate the overall performance (i.e., problem-solving accuracy) when generating responses at different target lengths. In our experiments, target lengths are selected from $\{512,1024,2048,3600\}$ tokens. For all our experiments, we run 16 random seeds, with a temperature of 0.6.
@@
Figure 2: Performance of L1 under varying token budgets. Pass rate (y-axis) across benchmarks as a function of tokens used (x-axis). L1-Exact and L1-Max significantly outperform length-controlled baseline (S1) while approaching performance of unconstrained models.
@@ 5 Results and Analysis
In this section, we report and analyze the effectiveness of the proposed method (LCPO) across various settings and benchmarks. We evaluate our method’s relative performance, generalization capability on out-of-domain tasks, controllability of length constraints and competitive performance in short CoT setups, and examine learned reasoning patterns.
@@
Figure 2 compares performance of L1-Exact and L1-Max with other baselines across varying generation lengths. Both variants of L1 achieve superior performance across all token budgets while maintaining strong length control. Compared to S1, the only other method specifically designed for length control, L1 shows remarkable improvements, over 100-150% relative and 20-25% absolute performance gains at both 512 and 1024 token budgets. This substantial difference can be attributed to two key factors: (1) L1 intelligently adapts its chain-of-thought to fit within specified length constraints without disrupting the reasoning process, while S1 often truncates mid-reasoning; and (2) L1 is explicitly trained to generate high-quality reasoning chains of varying lengths, effectively distilling reasoning patterns from longer chains to shorter ones.
@@
L1-Exact performs approximately 1% below DeepScaleR-4K, which is the same underlying model as L1, but trained without length constraints. However, this difference is primarily observed in the AIME dataset, where unconstrained models can generate very long chains for complex problems. Additionally, L1-Exact allocates the same token budget to all problems regardless of difficulty, potentially using extra tokens on simpler problems. L1-Max effectively alleviates this challenge, matching the performance of DeepScaleR-4K by optimizing token usage based on problem difficulty while respecting the upper ceiling. In doing so, it outperforms even L1-Exact often by up to 2x fewer tokens. L1-Max is particularly valuable when exact token counts are less desirable than a worst-case compute budget. Finally, the scaling trends suggest that with longer context training, L1 would match or even surpass DeepScaleR-24K’s performance while maintaining a high degree of length control.
@@
Figure 3 confirms that L1 generalizes robustly to new domains: performance consistently scales positively with token budget for OOD general reasoning datasets, approaching or matching DeepScaleR-4K benchmarks despite explicit length control constraints. For GPQA and LSAT, we observe the same linear performance scaling trend as in our primary datasets, with L1 matching DeepScaleR-4K’s performance at comparable token budgets. This generalization is particularly impressive given that L1 was not explicitly trained on these tasks. For MMLU, we see a less pronounced linear scaling relationship ($R^{2}=0.66$), likely because these knowledge-focused questions benefit less from extended reasoning.
@@
Figure 5: Budget Violation Rates across different datasets. L1-Max maintains less than 2.5% budget soft violation rate on across different token budgets, with lower violations at higher token lengths. On average the violation rate of mere 1.3%.
We quantitatively assess L1’s ability to follow length constraints across various mathematical reasoning datasets. As shown in Figure 5, our model maintains consistent control across all token budgets (512, 1024, 2048, and 3600 tokens), with observed output lengths usually closely matching the requested lengths. Further, in Figure 11, we show the mean error: ($\frac{E_{x\sim D}[n_{generated}]-n_{gold}}{n_{gold}}$), which captures the average deviation from target lengths across the dataset. The figure demonstrates that mean error is low: close to 3% for all math reasoning datasets. Although OOD datasets exhibit predictably higher errors (20-40%), these remain preferable over uncontrolled prompting. Further Analysis in Appendix A.2 demonstrates that larger errors primarily appear at higher token budgets on tasks like MMLU, where the longer chain of thoughts is mostly unnecessary. Additionally, in Appendix A.1, we show that error can be further reduced significantly with extended RL training.
Further, for L1-Max, we show the budget violation rates in Figure 5. In particular, we compute soft violation rates: $|n_{generated}-n_{gold}|>500$. The violation rates are low, ranging from 0.3% to 2.3% across different token budgets. The results demonstrate that L1-Max can be reliably used for ensuring length control.
@@
Table 1: Short CoT performance across math benchmarks. L1-Max surpasses its non-reasoning base model and competes with frontier models under identical short CoT lengths.
Given L1’s strong performance at lower token budgets, we conducted a focused evaluation comparing it to both its base non-reasoning model (Qwen-2.5-1.5B-Instruct) and significantly larger non-reasoning models (GPT-4o and Llama-3.3-70B) at comparable generation lengths. Table 1 presents these results, showing that L1 consistently outperforms or matches all models across all datasets despite using equivalent token budgets. Further, on average, L1 is 5% better than its non-reasoning counterpart, and even outperforms GPT-4o by 2% on average. We refer to these models as Short Reasoning Models (SRMs), given their shorter generation lengths, yet exhibiting similar patterns as full reasoning models (See Section 5).
@@
Figure 6: Keyword usage by reasoning pattern at 512 vs. 4096 tokens. L1 adaptively adjusts reasoning patterns based on requested token length.
Figure 7: Distribution of tokens for thinking vs. solution across chain-of-thought lengths.
To understand how L1 changes its reasoning approach across different length constraints, we analyzed how frequently certain reasoning-related terms appear in outputs of different lengths. Specifically, we calculated the normalized occurrence rate of most common reasoning terms in 512-token outputs compared to 4096-token outputs, showing how the model’s reasoning strategies shift when given different length constraints. Figure 6 organizes these keywords into 4 distinct reasoning patterns: “Self-Correction and Verification,” “Exploration and Alternatives,” “Context Setting,” and “Conclusion Drawing.”
Figure 6 shows that self-correction and verification keywords appear approximately twice as frequently in 4096-token outputs compared to 512-token outputs. Similarly, conclusion-drawing terms increase 2-10x with increased token budget, indicating more thorough solution validation and completion. Interestingly, most exploration-related keywords decrease in relative frequency at higher token counts, with “Alternatively” being a notable exception. Overall, we observe that smaller CoTs have reasoning patterns similar to their longer counterparts, but with changed relative frequencies that favor more self-verification and conclusion drawing in longer CoTs.
@@
Additional Experiments and Analysis
We further conduct several additional experiments and analysis to validate the robustness and generalizability of our approach across different configurations and model scales.
Reward Function Ablations. We systematically evaluate alternative reward function formulations for L1-Max in Appendix A.7 to understand the impact of different objective designs on both performance and length adherence. Our ablation study examines three variants: a single-objective approach using only the maximum length constraint, an additive formulation combining correctness and length penalties, and a sigmoid-based penalty function. While all variants achieve similar budget violation rates, the additive formulation exhibits undesirable behavior by collapsing to extremely short chain-of-thought sequences that trivially satisfy constraints but severely degrade reasoning quality. Our chosen multiplicative formulation in Equation 2 strikes an optimal balance, maintaining strong performance while ensuring robust adherence to length constraints.
Sampling Parameter Robustness. Our analysis in Appendix A.4 demonstrates that L1’s length control capabilities remain stable across varying decoding parameters. When evaluated across temperature settings from 0.0 to 1.0, the model maintains consistent performance with mean length deviations of only 5.6-6.3% from target constraints. This robustness is particularly important for practical deployment, as it ensures reliable length control regardless of the specific sampling strategy employed. The observed average token lengths (1765-1782) closely match the theoretical expectation of 1796 tokens when averaged across our standard evaluation lengths, indicating that the model’s learned length control generalize effectively beyond the sampling parameters used in RL training.
Parallel versus Sequential Scaling Trade-offs. In Appendix A.8, we investigate how L1 performs when computational budget is allocated through parallel sampling (majority voting) versus sequential scaling (longer chain-of-thought generation). Our findings reveal that sequential scaling consistently outperforms parallel approaches at equivalent token budgets, aligning with recent trends in test-time compute allocation. Nonetheless parallel scaling is still beneficial, as it may be desirable in cases where lower latency is desired.
Supervised Fine-tuning is not effective. Our analysis in Appendix A.6 highlight ineffectiveness of supervised fine-tuning for inducing length control in reasoning models. In particular, we generate multiple responses from the reasoning model, relabel reasoning chains with their token lengths and repurpose this generated data for training. Our results highlight that SFT-trained models consistently fail to follow length constraints, generating very long output regardless of requested lengths. We hypothesize this failure stems from two factors: the narrow distribution of output lengths for given questions leads models to ignore length instructions, and the absence of online length-sensitive rewards prevents acquisition of adaptive reasoning capabilities. The result highlights the effectiveness of reinforcement learning based LCPO for achieving high degree of length control.
Scaling to Larger Models. We demonstrate the scalability of our approach by applying LCPO to DeepSeek-R1-Distill-7B in Appendix A.9, showcasing the method’s potential for larger reasoning models. The 7B model exhibits identical trends to our 1.5B experiments: high length controllability, minimal budget violation rates, and strong performance relative to length-controlled baselines across token budgets.
@@ 6 Conclusion
In this work, we introduced Length Controlled Policy Optimization (LCPO), a simple yet powerful reinforcement learning-based method enabling adaptive control over the length of reasoning chains in language models. We use LCPO to train L1, a reasoning language model, optimized to produce outputs that adhere to length constraints given in its prompt. LCPO significantly surpasses previous test-time scaling methods, achieving over 100% relative and 20% absolute improvements in mathematical reasoning tasks compared to prior length-control approaches. Further, we demonstrated that L1 generalizes robustly beyond its training distribution, extending its length-controlled capabilities to out-of-domain tasks. Furthermore, our analysis revealed an intriguing phenomenon: models trained to generate longer reasoning chains become unexpectedly strong short reasoning models, outperforming significantly larger frontier models such as GPT-4o at identical generation lengths. By providing length control using simple prompt, LCPO opens promising avenues toward more efficient, flexible, and scalable reasoning models.
7 Limitations
Our results highlight strong length control capabilities of LCPO for both in-domain and out-of-domain tasks. However, we note that the models may not generalize to requested lengths that are longer than what they are trained for. It remains an important future direction to train models to use more inference compute at test than at training time. Further, we trained our models with length rewards on the complete outputs reflecting the real-world cost incurred by end-users. However, exploring other variants of rewards that, for instance, control the length of the reasoning tokens could be an interesting future direction.
8 Acknowledgements
We thank Convergent Research and the OpenAI Researcher Access program. This work was supported in part by the National Science Foundation under Grant Nos. DMS-2434614 and DMS-2502281. Pranjal is supported by SoftBank Group–Arm Fellowship.